Journalist Report – December 28th

Today, we had a lie-in and woke up at 8am. After the breakfast, we made sure to charge the generator, we played poker and prepared for the EVA. Ilaria, John and I then took soil samples on our ATVs and our rover. We hiked to the top of a hill and enjoyed the landscape. We were surrounded by a large canyon and desert plains. As we went back to the Hab, we did the last communication experiment of our mission. After this 3-hour-long EVA, we had a well-deserved lunch. Everybody went back to their daily activities: I cleaned the helmets and checked the electronic systems, David worked in the GreenLab, while Thibault, Ilaria and John did a medical experiment.


EVA Report – December 28th

Report: Crew 185 EVA #8 for SOL 10.

Propose of EVA:
-Collecting geological samples for soil density study. We will do multiple stops along Cow Dung road from the Hab to Gateway Lith. -We have done two stops along the way to collect samples, and it was a good teamwork!
-We stayed at the Lith for about 30 min for exploration and observing the panorama!
-Astro-CapCom communication. CapCom will drive the Astros to a specific geographical position (nearby the Hab) without the use of a GPS. EVA #8:
Start time: 1030am
End time: 1330pm->1430pm
Ilaria: Leader
Arno and John: Astros
David and Thibo are Comms.
Destination: Gateway to Lith (110.81 W – 38.45 N), see attachment. Vehicles: ATV (Arno), ATV (John), Rover Spirit (Ilaria)

Prepared by Ilaria Cinelli

GreenHab – December 28th

David Murray

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient with window/door open from 9:40 – 18:06
Working Hour: 15:38
Inside temp at working hour: 23 C
Outside temp during working hours: 8 C
Inside temperature high: 26 C
Inside temperature low: 15 C
Inside humidity: 60%

Functioning nominally
Turns on at 16 C and off at 20 C
Not functional
Shade Cloth On/Off
The greenhab lights are on from 17:00 to 24:00 every night
Average temperature:
Changes to Crops
Morning Research Observations
Daily Water Usage for Crops
6 gallons
Time(s) of Watering Crops

I replaced some of the tomato cages around the beans with wooden poles. Fourteen moringa seeds were planted in soil and five were started in damp paper towels.
Support/Supplies Needed

Sol Summary – December 28th

Sol 10

Long EVA

David Murray

Mission Status: Everything is going well

Sol Activity Summary: We woke up at 0800, had breakfast, and switched off the generator. Today, the EVA crew got a chance to gain experience on the ATV’s and to explore Lith Canyon. Then, they completed the navigation exercise and returned to the hab. After lunch, I planted Moringa oleifera seeds in the greenhab and watered the plants, Tibo and Ilaria participated in John’s nerve block experience, and Arno checked our state of charge. The crew plans to eat dinner and finish the rest of the movie that we started last night before going to bed.

Look Ahead Plan: Record our answers to questions from children around the world

Anomalies in work: None

Weather: Sunny, warm

Crew Physical Status: Healthy

EVA: Exploration, sample collection

Reports to be filed: Journalist Report, Greenhab Report, EVA Request, Sol Summary, EVA report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site